OpenStudy (anonymous):

x^2+8x+10=0 solve by completing the sq!

OpenStudy (bahrom7893):

x^2 + 2*4x + 16 - 6 = 0 (x^2+4)^2-6 = 0

OpenStudy (lalaly):

(x+4)^2-6=0 *

OpenStudy (bahrom7893):

woops sorry meant (x+4)^2-6=0

OpenStudy (bahrom7893):

(x+4-sqrt(6))(x+4+sqrt(6)) = 0

OpenStudy (bahrom7893):

x = sqrt(6)-4 x = -sqrt(6)-4 lala check lol

OpenStudy (anonymous):

tahnks so much!